# The Open Journal of Astrophysics

### v1

@@ Abstract
We present diffHOD-IA, a fully differentiable implementation of a halo occupation distribution (HOD) model that incorporates galaxy intrinsic alignments (IA). Motivated by the diffHOD framework of Horowitz et al. (2022), we create a new implementation that extends differentiable galaxy population modeling to include the orientation-dependent statistics crucial for weak gravitational lensing analyses. Our implementation combines this HOD formulation with the intrinsic alignment model of Van Alfen et al. (2024), enabling end-to-end automatic differentiation from HOD and IA parameters through to the galaxy field. We additionally extend this framework to differentiably model two-point correlation functions, including galaxy clustering and IA statistics. We validate diffHOD-IA against the reference halotools-IA implementation using the Bolshoi-Planck simulation, demonstrating excellent agreement across both 1-point and 2-point statistics. We verify the accuracy of gradients computed via automatic differentiation by comparing against finite difference estimates for both HOD and IA parameters. We present science use cases leveraging gradients in the simulations to uncover the IA parameters of a galaxy field representative of the tng300 simulation. We further use diffHOD-IA in a Hamiltonian Monte Carlo analysis and compare performance with halotools-IA and a neural-network based emulator of halotools-IA, named IAEmu. Unlike emulator-based approaches for statistics, diffHOD-IA provides differentiability at the catalog level, enabling its integration into field-level inference pipelines and extension to arbitrary summary statistics for next-generation weak lensing analyses.
@@ 1 Introduction
IA modeling has traditionally relied on analytic approaches rooted in perturbation theory, including the nonlinear alignment model (Hirata and Seljak, 2004; Bridle and King, 2007), the tidal alignment and tidal torquing (TATT) model (Blazek et al., 2015, 2019), and more recent effective field theory treatments (Vlah et al., 2020, 2021). While these analytic models provide valuable physical insight and computational efficiency, they often struggle to accurately capture nonlinear effects on small scales. Simulation-based approaches offer a complementary path forward. Fully nonlinear scales can be described with a halo model (Fortuna et al., 2021), which provides a framework for connecting galaxies to their host dark matter halos without explicit modeling of baryonic physics. This is in addition to semi analytic models, which combine perturbation theory approaches with halo-based methods (Maion et al., 2024). Hydrodynamic simulations provide direct predictions for galaxy alignments, although measurements exhibit significant variance across simulation suites depending on the sub-grid physics employed (Tenneti et al., 2016; Samuroff et al., 2021). For this reason, simulation frameworks that capture the galaxy-halo connection without strong dependence on subgrid physics can be valuable.
The halo occupation distribution (HOD) framework provides a powerful statistical description of the galaxy-halo connection (Peacock and Smith, 2000; Seljak, 2000; Berlind and Weinberg, 2002; Zheng and others, 2005; Zheng et al., 2007a). By specifying the probability that a halo of given mass hosts a certain number of galaxies, HOD models can populate dark matter halos from $N$-body simulations with realistic galaxy distributions. This approach has been widely successful in modeling galaxy clustering across a range of scales and redshifts (Zehavi and others, 2011; Coupon and others, 2012; Parejko and others, 2013). Extensions to the basic HOD framework have incorporated galaxy properties such as color, luminosity, and stellar mass (Zheng et al., 2007b; Zehavi and others, 2011), as well as secondary halo properties beyond mass (Hearin and others, 2016; Wechsler and Tinker, 2018).
@@
A key limitation of the HOD is its reliance on stochastic sampling procedures that are not differentiable, restricting Bayesian inference with HOD simulations to methods like Markov Chain Monte Carlo (MCMC). This precludes the use of gradient-based optimization and inference methods like Hamiltonian Monte Carlo (HMC; Duane et al. 1987; Neal 2011), which have proven highly efficient in machine learning and increasingly in cosmological applications (see Dvorkin et al. (2022) for a review). As Stage IV survey analyses begin to require inference over larger parameter spaces to account for systematics, MCMC-based inference becomes increasingly computationally demanding in high-dimensional settings. For instance, the Dark Energy Survey Year 6 result included inference over 50 nuisance parameters, in addition to cosmological parameters (Collaboration et al., 2026). Consequently, there is a growing need for more computationally efficient (differentiable) methods that enable tractable inference over many free parameters.
@@ 2.2 Satellite Occupation
We utilize a SubhaloPhaseSpace model for the IA in which satellite galaxies are always placed at the center of subhalos. These are preferentially placed in more massive subhalos first. In the event that the number of satellites exceeds the number of available subhalos, the satellite occupation resorts to a NFWPhaseSpace mode, where the remaining satellites are spatially distributed by sampling from a Navarro-Frenk-White (NFW) distribution (Navarro et al., 1996). One can also utilize the NFWPhaseSpace for all satellite occupation, in the event that subhalo information is not available.
@@ 2.3 Galaxy Intrinsic Alignments
$B(\kappa)=\frac{1}{2}\int_{0}^{1}e^{-\kappa t^{2}}\,dt.$ (6)
In this formulation, $\phi$ is sampled from a uniform distribution.
@@ 3 Differentiable halotools-IA
For galaxy clustering, our differentiable HOD methodology closely follows that of diffHOD (Horowitz et al., 2022), which we will summarize below. diffHOD was shown to agree with (non-differentiable) HOD-generated catalogs at the level of the power spectrum to within $\sim 2.5\%$ at the scales considered ($k<3\;h/\text{Mpc}$). We extend this framework by introducing a differentiable procedure for Dimroth–Watson sampling, employing inverse cumulative distribution function (CDF) sampling. As we adopt their methodology closely, we refrain from extensively benchmarking the HOD component of diffHOD-IA, and instead focus our analyses on the IA modeling and its accuracy compared to halotools-IA.
@@ 3.4 Differentiable NFW Sampling
In the event that $N_{\text{sat}}>|\mathcal{H}_{h}|$, we assign the locations of all remaining satellite galaxies according to samples from a NFW distribution. We generate satellite positions via an inverse CDF sampling procedure, which we will also use for IA sampling in Section 3.5. Our treatment follows the approach of Horowitz et al. (2022), but rather than using their closed-form approximation based on the Lambert $W$ function, we employ a numerically stable Newton iteration to invert the CDF. Both approaches are differentiable and similarly accurate. We begin with the CDF of the NFW profile:
@@
where $c$ is the concentration parameter of the host halo and $r_{\mathrm{vir}}$ is the virial radius of the host halo. We obtain position samples by drawing $u\sim\text{Uniform}(0,1)$ and solving for $r=P^{-1}(u)$ via Newton’s method for six iterations, which was found to be sufficient.
@@ 3.5 Differentiable Galaxy Intrinsic Alignments
To draw differentiable samples, we invert the CDF
$F(t\mid\kappa)=\frac{1}{Z(\kappa)}\int_{-1}^{t}\exp(-\kappa s^{2})\,ds,$ (21)
by solving for $t$ in $t=F^{-1}(u\mid\kappa)$, where $u\sim\mathrm{Uniform}(0,1)$. This yields the following piecewise expression for the CDF:
$F(t\mid\kappa)=\begin{cases}\frac{1}{2}\left[1+\dfrac{\operatorname{erf}(\sqrt{\kappa}t)}{\operatorname{erf}(\sqrt{\kappa})}\right],&\kappa>0\\ \frac{t+1}{2},&\kappa=0\\ \frac{1}{2}\left[1+\dfrac{\operatorname{erfi}(\sqrt{-\kappa}t)}{\operatorname{erfi}(\sqrt{-\kappa})}\right],&\kappa<0\end{cases}$ (22)
As we are interested in solving $t=F^{-1}(u\;|\;\kappa)$, we must invert the expression in Equation 22 for $t$. To simplify notation, we apply the transformation $u^{\prime}=2u-1$, which maps $u\sim\mathrm{Uniform}(0,1)$ to $u^{\prime}\sim\mathrm{Uniform}(-1,1)$, and drop the prime for clarity. This yields:
@@
In practice, the inverse CDF is solved via Newton’s method with six iterations. The resulting galaxy orientation vectors are fully differentiable with respect to the IA parameters.
@@ 4 Differentiable Correlation Functions
Typical cosmological analyses employ low dimensional summary statistics computed over the full galaxy field. We proceed by outlining a prescription to differentiably calculate 2PCFs, enabling an end-to-end differentiable pipeline from halo occupation and IA parameters to summary statistics. We find that computing the IA statistics is directly differentiable with respect to the IA parameters using discrete galaxy catalogs; computing $\xi(r)$ requires generating catalogs where the galaxies have a weight proportional to their occupation probability, as done in Hearin et al. (2022). We note that the IA statistics $\omega(r)$ and $\eta(r)$ also depend on the HOD parameters through the galaxy positions and number counts; however, for the experiments in this work, we operate in a fixed HOD setting to validate the differentiable IA implementation. We develop and benchmark accuracy for both the weighted $\xi(r)$ and IA correlation estimators in Appendix B. All correlation measurements use 20 logarithmically-spaced bins from $r=0.1\;h^{-1}\text{Mpc}$ to $r=16\;h^{-1}\text{Mpc}$.
@@ 4.1 Intrinsic Alignment Statistics
In this context, there are two IA statistics of interest: $\omega(r)$ and $\eta(r)$. The galaxy position-orientation correlation function $\omega(r)$ is defined as:
@@
where $\hat{e}$ is the galaxy orientation unit vector that specifies the intrinsic orientation of each galaxy’s major axis, $\mathbf{x}$ is the position vector of a galaxy, and $\hat{r}$ is the unit vector in the direction of galaxy separation. The galaxy orientation-orientation correlation function $\eta(r)$ is defined as:
@@
which measures the correlation between the orientations of galaxy pairs as a function of their separation.
@@ 5.1 Comparison to halotools-IA
In the bottom panels of Figure 1, we compare $\xi(r)$, $\omega(r)$, and $\eta(r)$ between the two implementations. This is using the (non-differentiable) halotools-IA implementation for the correlation estimators, to highlight any potential differences at the 2PCF level coming solely from the galaxy occupation and IA. We find generally excellent agreement between diffHOD-IA and halotools-IA across all three statistics. This even includes the IA statistic $\eta(r)$, which is much noisier than $\omega(r)$ due to galaxy shape noise. We additionally see comparable error bars between the implementations across 100 realizations. We benchmark the accuracy of the differentiable correlation estimators as outlined in Section 4 in Appendix B.
@@ 6 Applications of differentiable capabilities
We demonstrate the utility of diffHOD-IA with gradient-based optimization experiments. These experiments illustrate how gradients in the simulation can be used to obtain parameter estimates and posteriors over parameters of interest from mock observational data. All experimental results presented use a value of $N_{\text{max}}=48$, $\tau=0.1$, and SubhaloPhaseSpace with a NFWPhaseSpace fallback. A constant radial alignment strength model is used for the subhalo IA. As our construction follows Horowitz et al. (2022) closely, and having validated both the accuracy and differentiability of diffHOD-IA against halotools-IA, our emphasis will be on the IA parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$; several experiments illustrating the differentiability of the HOD are included in Horowitz et al. (2022). We construct moment-matching and differentiable correlation function optimization objectives. In addition, we demonstrate accelerated inference with diffHOD-IA using HMC, compared to halotools-IA with MCMC.
@@ 6.1 Moment-matching Objective
Figure 4: Gradient-based recovery of IA parameters from 50 random initializations using moment-matching optimization. The target parameters ($\mu_{\rm cen}=0.7905$, $\mu_{\rm sat}=0.307$, pink star) represent the best-fit $\mu$ values to the tng300 HOD configuration, with the empirical uncertainty shown as pink contours. Optimization trajectories are shown as gray lines connecting initial positions to converged solutions. Brown circles denote optimizations using a single HOD realization seed per gradient step, while green squares show results when averaging over three seeds. The inset panel (lower left) shows a zoomed view of the convergence region, revealing tight clustering of final parameter estimates around the true values. Both single-seed and three-seed strategies successfully recover the target parameters across diverse initializations.

### v2

@@ Abstract
We present diffHOD-IA, a fully differentiable implementation of a halo occupation distribution (HOD) model that incorporates galaxy intrinsic alignments (IA). Motivated by the diffHOD framework of Horowitz et al. (2022), we create a new implementation that extends differentiable galaxy population modeling to include the orientation-dependent statistics crucial for weak gravitational lensing analyses. Our implementation combines this HOD formulation with the intrinsic alignment model of Van Alfen et al. (2024), enabling end-to-end automatic differentiation from HOD and IA parameters through to the galaxy field. We additionally extend this framework to differentiably model two-point correlation functions, including galaxy clustering and IA statistics. We validate diffHOD-IA against the reference halotools-IA implementation using the Bolshoi-Planck simulation, demonstrating agreement within sample variance in galaxy number counts and at $2\%$ error in two-point statistics. We verify the accuracy of gradients computed via automatic differentiation by comparing against finite difference estimates for both HOD and IA parameters. We present science use cases that leverage gradients in the parameter space to find the IA of a galaxy field representative of the tng300 simulation. We further use diffHOD-IA in a Hamiltonian Monte Carlo analysis and compare performance with halotools-IA and a neural-network based emulator of halotools-IA, named IAEmu. Unlike emulator-based approaches for statistics, diffHOD-IA provides differentiability at the catalog level, enabling its integration into field-level inference pipelines and extension to arbitrary summary statistics for next-generation weak lensing analyses.
@@ 1 Introduction
IA modeling has traditionally relied on analytic approaches rooted in perturbation theory, including the nonlinear alignment model (Hirata and Seljak, 2004; Bridle and King, 2007), the tidal alignment and tidal torquing (TATT) model (Blazek et al., 2015, 2019), and more recent effective field theory treatments (Vlah et al., 2020, 2021). While these analytic models provide valuable physical insight and computational efficiency, they often struggle to accurately capture nonlinear effects on small scales. Simulation-based approaches offer a complementary path forward. Fully nonlinear scales can be described with a halo model (Fortuna et al., 2021), which provides a framework for connecting galaxies to their host dark matter halos without explicit modeling of baryonic physics. This is in addition to semi-analytic models, which combine perturbation theory approaches with halo-based methods (Maion et al., 2024). Hydrodynamic simulations provide direct predictions for galaxy alignments, although measurements exhibit significant variance across simulation suites depending on the sub-grid physics employed (Tenneti et al., 2016; Samuroff et al., 2021; van Heukelum and Chisari, 2026). For this reason, simulation frameworks that capture the galaxy-halo connection without strong dependence on subgrid physics can be valuable.
The halo occupation distribution (HOD) framework provides a powerful statistical description of the galaxy-halo connection (Peacock and Smith, 2000; Seljak, 2000; Berlind and Weinberg, 2002; Zheng and others, 2005; Zheng et al., 2007a). By specifying the probability that a halo of given mass hosts a certain number of galaxies, HOD models can populate dark matter halos from $N$-body simulations with realistic galaxy distributions. This approach has been widely successful in modeling galaxy clustering across a range of scales and redshifts (Zehavi and others, 2011; Coupon and others, 2012; Parejko and others, 2013). Extensions to the basic HOD framework have incorporated galaxy properties such as color, luminosity, and stellar mass (Zheng et al., 2007b; Zehavi and others, 2011), as well as secondary halo properties beyond mass (Hearin and others, 2016; Wechsler and Tinker, 2018). Importantly, a growing body of observational evidence from photometric, spectroscopic, and narrowband surveys has established that IA is strongly correlated with these same galaxy properties (Johnston et al., 2019; Joachimi and others, 2013b; Fortuna et al., 2021, 2024; Samuroff et al., 2023; Georgiou et al., 2025; Siegel et al., 2025; Navarro-Gironés et al., 2025), suggesting that these HOD frameworks are naturally well-suited to model these dependencies compared to analytic approaches such as NLA and TATT.
@@
A key limitation of the HOD is its reliance on stochastic sampling procedures that are not differentiable, restricting Bayesian inference with HOD simulations to methods like Markov Chain Monte Carlo (MCMC). This precludes the use of gradient-based optimization and inference methods like Hamiltonian Monte Carlo (HMC; Duane et al. 1987; Neal 2011), which have proven highly efficient in machine learning and increasingly in cosmological applications (see Dvorkin et al. (2022) for a review). As Stage IV survey analyses begin to require inference over larger parameter spaces to account for systematics, MCMC-based inference becomes increasingly computationally demanding in high-dimensional settings. For instance, the Dark Energy Survey Year 6 result included inference over 50 nuisance parameters, in addition to cosmological parameters (Collaboration et al., 2026). In Piras et al. (2024), it was estimated that a Stage IV full 3x2 analysis would require up to 12 years of compute time on 48 CPU cores with MCMC-like sampling, which is contrasted with 8 days on GPU with differentiable sampling techniques. Consequently, there is a growing need for more computationally efficient (differentiable) methods that enable tractable inference over many free parameters.
@@ 2.2 Satellite Occupation
We utilize a SubhaloPhaseSpace model for the IA in which satellite galaxies are always placed at the center of subhalos. These are preferentially placed in more massive subhalos first. In the event that the number of satellites exceeds the number of available subhalos, the satellite occupation resorts to a NFWPhaseSpace mode, where the remaining satellites are spatially distributed by sampling from a Navarro-Frenk-White (NFW) distribution (Navarro et al., 1996) about the central halo. One can also utilize the NFWPhaseSpace for all satellite occupation, in the event that subhalo information is not available.
@@ 2.3 Galaxy Intrinsic Alignments
$B(\kappa)=\frac{1}{2}\int_{0}^{1}e^{-\kappa t^{2}}\,dt,$ (6)
where $t=\cos(\theta)$. In this formulation, $\phi$ is sampled from a uniform distribution.
@@ 3 Differentiable halotools-IA
For galaxy clustering, our differentiable HOD methodology closely follows that of diffHOD (Horowitz et al., 2022), which we will summarize below. diffHOD was shown to agree with (non-differentiable) HOD-generated catalogs at the level of the power spectrum to within $1\%$ at the scales considered ($k<3\;h/\text{Mpc}$), which is generally sufficient for Stage IV cosmological analyses. We extend this framework by introducing a differentiable procedure for Dimroth–Watson sampling, employing inverse cumulative distribution function (CDF) sampling. As we adopt their methodology closely, we refrain from extensively benchmarking the HOD component of diffHOD-IA, and instead focus our analyses on the IA modeling and its accuracy compared to halotools-IA.
@@ 3.4 Differentiable NFW Sampling
In the event that $N_{\text{sat}}>|\mathcal{H}_{h}|$, we assign the locations of all remaining satellite galaxies according to samples from an NFW distribution. We generate satellite positions via an inverse CDF sampling procedure, which we will also use for IA sampling in Section 3.5. Our treatment follows the approach of Horowitz et al. (2022), but rather than using their closed-form approximation based on the Lambert $W$ function, we employ a numerically stable Newton iteration to invert the CDF. Both approaches are differentiable and similarly accurate. We begin with the CDF of the NFW profile:
@@
where $c$ is the concentration parameter of the host halo and $r_{\mathrm{vir}}$ is the virial radius of the host halo. We obtain position samples by drawing $u\sim\text{Uniform}(0,1)$ and solving for $r=P^{-1}(u)$ via Newton’s method for six iterations, which was found to be sufficient yielding a convergence error of $\lesssim 10^{-4}$.
@@ 3.5 Differentiable Galaxy Intrinsic Alignments
To draw differentiable samples, we define the CDF
$F(t\mid\kappa)=\frac{1}{Z(\kappa)}\int_{-1}^{t}\exp(-\kappa s^{2})\,ds.$ (21)
This integral admits closed-form expressions that depend on the sign of $\kappa$, yielding the following piecewise form of the CDF:
$F(t\mid\kappa)=\begin{cases}\frac{1}{2}\left[1+\dfrac{\operatorname{erf}(\sqrt{\kappa}t)}{\operatorname{erf}(\sqrt{\kappa})}\right],&\kappa>0\\ \frac{t+1}{2},&\kappa=0\\ \frac{1}{2}\left[1+\dfrac{\operatorname{erfi}(\sqrt{-\kappa}t)}{\operatorname{erfi}(\sqrt{-\kappa})}\right],&\kappa<0.\end{cases}$ (22)
To generate samples, we invert the CDF by solving $t=F^{-1}(u\mid\kappa)$ for $u\sim\mathrm{Uniform}(0,1)$. For convenience, we rescale the uniform random variable via the transformation $u^{\prime}=2u-1$, which maps $u\in(0,1)$ to $u^{\prime}\in(-1,1)$; we drop the prime in what follows. Inverting the appropriate branch of Equation 22 then yields:
@@
While Equation 23 admits analytic inverse forms, the implementation evaluates the inverse CDF analytically when possible, and otherwise solves the inverse relation numerically using a small number of Newton iterations when the inverse involves $\text{erfi}^{-1}$, which is not available in standard libraries. The resulting galaxy orientation vectors are fully differentiable with respect to the IA parameters.
@@ 4 Differentiable Correlation Functions
Typical cosmological analyses employ low dimensional summary statistics computed over the full galaxy field. We proceed by outlining a prescription to differentiably calculate 2PCFs, enabling an end-to-end differentiable pipeline from halo occupation and IA parameters to summary statistics. We find that computing the IA statistics is directly differentiable with respect to the IA parameters using discrete galaxy catalogs; computing $\xi(r)$ requires generating catalogs where the galaxies have a weight proportional to their occupation probability, as done in Hearin et al. (2022). We note that the IA statistics $\omega(r)$ and $\eta(r)$ also depend on the HOD parameters through the galaxy positions and number counts; however, for the experiments in this work, we operate in a fixed HOD setting to validate the differentiable IA implementation. We develop and benchmark the accuracy for both the weighted $\xi(r)$ and IA correlation estimators in Appendix B. All correlation measurements use 20 logarithmically-spaced bins from $r=0.1\;h^{-1}\text{Mpc}$ to $r=16\;h^{-1}\text{Mpc}$. These scales capture the transition from the 1-halo to 2-halo regime, while also allowing correlations to still be measured quickly.
@@ 4.1 Intrinsic Alignment Statistics
In this context, there are two IA statistics of interest: $\omega(r)$ and $\eta(r)$. Our treatment does not currently include full galaxy shape information, and the correlations as defined below are only sensitive to galaxy orientations. The galaxy position-orientation correlation function $\omega(r)$ is defined as:
@@
where $\hat{e}$ is the galaxy orientation unit vector that specifies the intrinsic orientation of each galaxy’s major axis, $\mathbf{x}$ is the position vector of a galaxy, and $\hat{r}$ is the unit vector in the direction of galaxy separation in 3D. The galaxy orientation-orientation correlation function $\eta(r)$ is defined as:
@@
which measures the correlation between the orientations of galaxy pairs as a function of their separation. The factor of $1/3$ in these equations account for the fact that
$\frac{1}{4\pi}\int_{0}^{2\pi}\int_{0}^{\pi}\cos^{2}\theta\sin\theta\;\text{d}\theta\;\text{d}\phi=\frac{1}{3},$ (27)
where integrating $\cos^{2}(\theta)$ over the sphere corresponds to the case of random alignments.
@@ 5.1 Comparison to halotools-IA
In the bottom panels of Figure 1, we compare $\xi(r)$, $\omega(r)$, and $\eta(r)$ between the two implementations. This is using the (non-differentiable) halotools-IA implementation for the correlation estimators, to highlight any potential differences at the 2PCF level coming solely from the galaxy occupation and IA. We find generally excellent visual agreement between diffHOD-IA and halotools-IA across all three statistics. This even includes the IA statistic $\eta(r)$, which is much noisier than $\omega(r)$ due to galaxy shape noise. We additionally see comparable error bars between the implementations across 100 realizations. This is confirmed upon inspecting the fractional error of the correlations, from which we find a mean bias of $0.28\%$ for $\xi(r)$, $0.16\%$ for $\omega(r)$, and $1.77\%$ for $\eta(r)$. We benchmark the accuracy of the differentiable correlation estimators as outlined in Section 4 in Appendix B.
@@ 6 Applications of differentiable capabilities
We demonstrate the utility of diffHOD-IA with gradient-based optimization experiments. These experiments illustrate how gradients in the simulation can be used to obtain parameter estimates and posteriors over parameters of interest from mock observational data. All experimental results presented use a value of $N_{\text{max}}=48$, $\tau=0.1$, and SubhaloPhaseSpace with a NFWPhaseSpace fallback. A constant radial alignment strength model is used for the subhalo IA. This combination of phase space and alignment model is sufficient to accurately model the IA of tng300, as studied in Van Alfen et al. (2024). As our construction follows Horowitz et al. (2022) closely, and having validated both the accuracy and differentiability of diffHOD-IA against halotools-IA, our emphasis will be on the IA parameters $\mu_{\text{cen}}$ and $\mu_{\text{sat}}$; several experiments illustrating the differentiability of the HOD are included in Horowitz et al. (2022). We construct moment-matching and differentiable correlation function optimization objectives. In addition, we demonstrate accelerated inference with diffHOD-IA using HMC, compared to halotools-IA with MCMC.
@@ 6.1 Moment-matching Objective
Figure 4: Gradient-based recovery of IA parameters from 50 random initializations using moment-matching optimization. The target parameters ($\mu_{\rm cen}=0.7905$, $\mu_{\rm sat}=0.307$, pink star) represent the best-fit $\mu$ values to the tng300 HOD configuration, with uncertainty contours in pink estimated from the scatter in effective $\mu$ values recovered by matching $\langle t^{2}\rangle$ to the theoretical Dimroth–Watson expectation across 50 independent HOD realizations. Optimization trajectories are shown as gray lines connecting initial positions to converged solutions. Gray circles denote optimizations using a single HOD realization seed per gradient step, while green squares show results when averaging over three seeds. The inset panel (lower left) shows a zoomed view of the convergence region, revealing tight clustering of final parameter estimates around the true values. Both single-seed and three-seed strategies successfully recover the target parameters across diverse initializations.
@@
We note that the roughly one order of magnitude larger uncertainty for $\mu_{\text{sat}}$ compared to $\mu_{\text{cen}}$ is primarily due to the smaller alignment strength of $\mu_{\text{sat}}=0.30$ relative to $\mu_{\text{cen}}=0.79$ At lower $\mu$, the Dimroth–Watson distribution is less sharply peaked and galaxy orientations are more dispersed, producing greater variance in $\mu$ inferred from a finite catalog. This larger uncertainty is reflected in the scatter of converged $\mu_{\text{sat}}$ values seen in Figure 4.